OpenStudy (anonymous):

a and b together can do a piece of work in 12 days. b and c together can do it in 15 days. if a is twice as good as c, in how many days a alone will do the same work?

OpenStudy (anonymous):

it should be 30 days

OpenStudy (anonymous):

how do you solve this can you explain...

OpenStudy (anonymous):

let the work done by a,b and c in one day be; A,B and C , respectively 1st case: a & b work done = 12A+12B 2nd case: b & c work done= 15B+15C now work done in both case is same so, 12A+12B=15B+15C => 12A=3B+15C => 4A= B+5C......(i) also it is given that a is twice as good as c so A=2C put this in eq(i) => 8C= B+5C => B=3C now again writing the work done 15B+15C= 15x3C + 15C= 60 C= 30A that means A alone can do the work in 30 days i hope the explanation is clearly understood